# Cybersecurity of Quantum Key Distribution Implementations

### v1

@@ 6.3 Examples
We will now analyze several examples of attacks on QKD implementations under our new definition.
@@ 6.3.1 Quantum Side-Channel Attacks
6.3.2 Non-Side-Channel Attacks
@@
This attack, discussed in Section 3.2.1, utilizes a vulnerability in Alice’s transmitter, which makes Alice’s sent states sometimes include more than one photon. Eve checks the number of photons in a pulse, and if there is more than one, she saves it as her private ancilla and sends the rest to Bob. Eve measures her saved photon after the bases are published, gaining perfect knowledge of the secret bit. Since the attack only utilizes Alice’s signal and generates states that Bob can measure, it is not a Quantum Side-Channel Attack.
@@
This attack works against receivers with different detectors for different bit values, where the detection sensitivities of each detector are unequal. Eve selectively delays Alice’s signal, independently of Alice’s choice of state, to make one bit value much more likely in Bob’s detection. The input of the attack is exactly Alice’s signal, and the output of the attack is time-shifted states, inside the space of states that Bob’s device can measure. As such, the Time-Shift attack is not a Quantum Side-Channel Attack.
@@
This attack was deliberately constructed to affect all of (and only) the extended Hilbert space that affects Bob’s measurement, and it does not depend on any input other than Alice’s signal. As such, this attack too is not a Quantum Side-Channel Attack.
Further classification of attacks under the Quantum Side-Channel Attack definition can be found in Section 8.
@@ 7.1 The Bright Illumination Attack
In the QKD system targeted by [56], the implemented protocol is polarization-based BB84, and measurement is performed using a passive basis choice and four detectors, as shown in Figure 1a.
@@
In the attack, Eve sends high-intensity light pulses to Bob’s device before the intended detection window. This makes the APDs operate in linear mode during the detection window, making them blind to low-intensity pulses. Once the APDs are in linear mode, Eve performs the “measure-resend” attack described in Figure 1b: she picks a specific measurement basis, measures Alice’s signal in that basis, and then generates a signal in the same polarization whose intensity is slightly above $2P_{th}$, where $P_{th}$ is defined as the minimal intensity the APDs can detect in linear mode.
When the pulse enters the polarization-independent beam splitter in Bob’s device, it is split into two pulses with intensity slightly higher than $P_{th}$. In the sub-component that measures in the basis Eve chose, the polarizing beam splitter will transfer the entire pulse to one arm, which will cause a click in the intended detector. In the other sub-component, the polarizing beam splitter will split the pulse into two parts with intensity around $P_{th}/2$, which will not cause a click in either detector.
@@ 8.1 Relationship to Quantum Side-Channel Attacks
There are Quantum Side-Channel Attacks against Alice and against Bob, but non-side-channel attacks also exist for both.
The diagram, as well as Section 6.3, show that Quantum Side-Channel Attacks are not only a theoretical definition: there are several practical examples of Quantum Side-Channel Attacks, both against Alice and against Bob, such as the Large Pulse attacks, the Injection Locking attack, and the Bright Illumination attacks.
However, in contrast to earlier definitions of side-channel attacks in QKD (see, e.g.,  [55]), our definition does not capture all implementation attacks: many (if not most) well-known examples do not use elements that are external to Alice and Bob’s protocol, and thus do not fall under our definition. Examples of such attacks are the Photon-Number-Splitting attack, the Trojan Pony attack, and the Time-Shift attack. Hence, our definition of Quantum Side-Channel Attacks represents a class of special, yet practical implementation attacks.
@@ 8.3 Relationship to Quantum Fuzzing
Quantum Fuzzing can discover both Quantum Side-Channel Attacks and non-side-channel attacks, but cannot discover all attacks against Bob. Furthermore, our current strategy cannot be applied against Alice’s device.
Since Quantum Fuzzing tests the behavior of a device by sending different input states to it, it can reveal the vulnerabilities behind several attacks. Four attacks that are triggered by sending specific input states to Bob’s device are the Large Pulse attack (against Bob), the Bright Illumination attacks, the Detector Efficiency Mismatch attacks, and the Time-Shift attack. The first two attacks are examples of Quantum Side-Channel Attacks, and the last two attacks are not Quantum Side-Channel Attacks.
@@ 9 Conclusions and Future Work
Expanding on the connection between classical cybersecurity and QKD implementations, we have presented three additional contributions. First, through our definition of Quantum Fuzzing (Section 4), which is the first vulnerability research method for QKD implementations, QKD system designers can search for issues in their implementation that they did not know existed. Second, our Reversed-Space Attacks methodology (Section 5) is a new generic exploit which computes and utilizes the QKD receiver’s attack surface, allowing system designers and attackers to build more sophisticated implementation attacks. Third, our definition of “Quantum Side-Channel Attacks” (Section 6) accounts for the special nature of side channels in QKD devices, and emphasizes that the security of QKD devices can be compromised by elements outside of Alice’s transmission and Bob’s measurement.
@@ Appendix E Formal Definition of Quantum Side-Channel Attacks
This is the formal formulation of Definition 8, presented in Section 6.
@@
  • Let $\mathcal{H}^{\mathrm{A}}\subseteq\mathcal{H}^{\mathrm{A}}_{\mathrm{full}}$ be the span of Alice’s sent states, and let $\mathcal{H}^{\mathrm{B}}\subseteq\mathcal{H}^{\mathrm{B}}_{\mathrm{full}}$ be the space of states that affect Bob’s measurement [^10] Note that in Section 5, this space is denoted by $\mathcal{H}^{\mathrm{P}}$..
@@
Eve’s attack is a Quantum Side-Channel Attack if either of these two conditions is true:
@@
    Eve’s attack depends non-trivially on $\mathcal{H}^{\mathrm{env}}$,
@@
    or the output of Eve’s attack non-trivially affects either $\mathcal{H}^{\mathrm{B}}_{\mathrm{full}}\setminus\mathcal{H}^{\mathrm{B}}$ or $\mathcal{H}^{\mathrm{env}}$.
@@ G.1 List of Individual Attacks
    This attack is discussed in Section 2.5.1 of [30].
  • Injection-Locking Attack [22].
@@
    This attack is applicable in QKD receivers that use different photo-detectors to detect different states, and the detection windows of the detectors are not identical. Eve selectively makes Alice’s signal arrive earlier or later than the original detection time, making only one detection result possible. This attack does not rely on Eve measuring Alice’s signal to generate an appropriate state to send to Bob, which makes it highly practical.
@@ G.2 List of Attack Families
    This is a general family of intercept-resend attacks where Eve sends malicious signals to Bob, which cause him to get a result of her choice. Details appear in Section 2.5.2.1 of [30].

### v2

@@ 6.3 Examples
We will now analyze several examples of attacks on QKD implementations under our new definitions.
@@ 6.3.1 Quantum Side-Channel Attacks
6.3.2 Quantum State-Channel Attacks
@@
This attack, discussed in Section 3.2.1, utilizes a vulnerability in Alice’s transmitter, which makes Alice’s sent states sometimes include more than one photon. Eve checks the number of photons in a pulse, and if there is more than one, she saves it as her private ancilla and sends the rest to Bob. Eve measures her saved photon after the bases are published, gaining perfect knowledge of the secret bit. Since the attack only utilizes Alice’s signal and generates states that Bob can measure, it is a Quantum State-Channel Attack.
@@
This attack works against receivers with different detectors for different bit values, where the detection sensitivities of each detector are unequal. Eve selectively delays Alice’s signal, independently of Alice’s choice of state, to make one bit value much more likely in Bob’s detection. The input of the attack is exactly Alice’s signal, and the output of the attack is time-shifted states, inside the space of states that Bob’s device can measure. As such, the Time-Shift attack is a Quantum State-Channel Attack.
@@
This attack was deliberately constructed to affect all of (and only) the extended Hilbert space that affects Bob’s measurements, and it does not depend on any input other than Alice’s signal. As such, this attack too is a Quantum State-Channel Attack.
Trojan-Pony Attack [55]:
In this attack, Bob’s vulnerability is that his detectors cannot count the number of photons that arrive in a pulse, and he interprets all “double-click” events (where more than one detector clicks) as signal loss, and not as an invalid result. Eve can use this vulnerability to launch a faked-states attack, sending states that Bob will either measure as a loss or as Eve’s choice of state. Since the attack assumes Alice is ideal and only affects the space Bob measures, it is a Quantum State-Channel Attack.
Imperfect Faraday Mirror Attack [68]:
In this attack, a fault in Alice’s optical setup makes her states span a three-dimensional Hilbert space instead of the standard two-dimensional qubit space. Eve performs a measure-resend attack, achieving a stronger distinction between the four states in the three-dimensional space, gaining significant information on the key while keeping the error rate below the required threshold. Since the attack relies on Alice’s sent states (and outputs valid qubits), it is also a Quantum State-Channel Attack.
Further classification of attacks under the Quantum Side-Channel Attack and Quantum State-Channel Attack definitions can be found in Section 8.
@@ 7.1 The Bright Illumination Attack
In the QKD system targeted by [56], the implemented protocol is polarization-based BB84, and measurement is performed using a passive basis choice and four detectors, as shown in Figure 1 (a).
@@
In the attack, Eve sends high-intensity light pulses to Bob’s device before the intended detection window. This makes the APDs operate in linear mode during the detection window, making them blind to low-intensity pulses. Once the APDs are in linear mode, Eve performs the “measure-resend” attack described in Figure 1 (b): she picks a specific measurement basis, measures Alice’s signal in that basis, and then generates a signal in the same polarization whose intensity is slightly above $2P_{\mathrm{th}}$, where $P_{\mathrm{th}}$ is defined as the minimal intensity the APDs can detect in linear mode.
When the pulse enters the polarization-independent beam splitter in Bob’s device, it is split into two pulses with intensity slightly higher than $P_{\mathrm{th}}$. In the sub-component that measures in the basis Eve chose, the polarizing beam splitter will transfer the entire pulse to one arm, which will cause a click in the intended detector. In the other sub-component, the polarizing beam splitter will split the pulse into two parts with intensity around $P_{\mathrm{th}}/2$, which will not cause a click in either detector.
@@ 8.1 Relationship to Quantum Side-Channel Attacks
There are Quantum Side-Channel Attacks and Quantum State-Channel Attacks against Alice and against Bob.
The diagram, as well as Section 6.3, shows that Quantum Side-Channel Attacks are not only a theoretical definition: there are several practical examples of such attacks both against Alice and against Bob, including the Large Pulse attacks, the Injection Locking attack, and the Bright Illumination attack family.
However, in contrast to earlier definitions of side-channel attacks in QKD (see, e.g., [55]), many practical attacks are not Quantum Side-Channel Attacks: many (if not most) well-known attack examples do not use elements that are external to Alice and Bob’s realistic protocol, and are therefore not Quantum Side-Channel Attacks, but rather Quantum State-Channel Attacks. Examples of such attacks are the Photon-Number-Splitting attack, the Trojan Pony attack, and the Time-Shift attack. Hence, our definitions of Quantum Side-Channel Attacks and Quantum State-Channel Attacks represent a meaningful distinction between implementation attacks.
@@ 8.3 Relationship to Quantum Fuzzing
Quantum Fuzzing can discover both Quantum Side-Channel Attacks and Quantum State-Channel Attacks, but cannot discover all attacks against Bob. Furthermore, our current strategy cannot be applied against Alice’s device.
Since Quantum Fuzzing tests the behavior of a device by sending different input states to it, it can reveal the vulnerabilities behind several attacks. Four attacks that are triggered by sending specific input states to Bob’s device are the Large Pulse attack (against Bob), the Bright Illumination attacks, the Detector Efficiency Mismatch attacks, and the Time-Shift attack. The first two attacks are examples of Quantum Side-Channel Attacks, and the last two attacks are Quantum State-Channel Attacks.
@@ 9 Conclusions and Future Work
Expanding on the connection between classical cybersecurity and QKD implementations, we have presented three additional contributions. First, through our definition of Quantum Fuzzing (Section 4), which is the first vulnerability research method for QKD implementations, QKD system designers can search for issues in their implementation that they did not know existed. Second, our Reversed-Space Attacks methodology (Section 5) is a new generic exploit which computes and utilizes the QKD receiver’s attack surface, allowing system designers and attackers to build more sophisticated implementation attacks. Third, our definitions of “Quantum Side-Channel Attacks” and “Quantum State-Channel Attacks” (Section 6) account for the special nature of side channels in QKD devices, and emphasize that the security of QKD devices can be compromised by elements both inside and outside of Alice’s transmission and Bob’s measurement.
@@ Appendix E Formal Definition of Quantum Side-Channel Attacks
This is the formal formulation of Definitions 8 and 9, presented in Section 6.
@@
  • Let $\mathcal{H}^{\mathrm{A}}\subseteq\mathcal{H}^{\mathrm{A}}_{\mathrm{full}}$ be the span of Alice’s sent states, and let $\mathcal{H}^{\mathrm{B}}\subseteq\mathcal{H}^{\mathrm{B}}_{\mathrm{full}}$ be the space of states that affect Bob’s measurement.
@@
Eve’s attack is a Quantum Side-Channel Attack if either of the following two conditions is true:
@@
    Eve’s attack depends non-trivially on $\mathcal{H}^{\mathrm{env}}$;
@@
    the output of Eve’s attack non-trivially affects either $\mathcal{H}^{\mathrm{B}}_{\mathrm{full}}\setminus\mathcal{H}^{\mathrm{B}}$ or $\mathcal{H}^{\mathrm{env}}$.
Definition 11.
Eve’s attack is a Quantum State-Channel Attack if the following two conditions are both true:
  1. 1.
    Eve’s attack does not depend non-trivially on $\mathcal{H}^{\mathrm{env}}$;
  2. 2.
    the output of Eve’s attack non-trivially affects only $\mathcal{H}^{\mathrm{B}}$ and $\mathcal{H}^{\mathrm{anc}}_{\mathrm{out}}$.
@@ G.1 List of Individual Attacks
    This attack is discussed in Section 3.2.1 of our work. Additional details appear in Section 2.5.1 of [30].
  • Injection-Locking Attack [22]
@@
    This attack is applicable in QKD receivers that use different photo-detectors to detect different states, and the detection windows of the detectors are not identical. Eve selectively makes Alice’s signal arrive earlier or later than the original detection time, making only one detection result possible. This attack does not rely on Eve measuring Alice’s signal to generate an appropriate state to send to Bob, which makes the attack highly practical.
@@ G.2 List of Attack Families
    This is a general family of intercept-resend attacks where Eve sends malicious signals to Bob, which cause him to get a result of her choice, as discussed in Section 3.2.3. Additional details appear in Section 2.5.2.1 of [30].